OpenStudy (anonymous):

20/15-x=5/x how do i solve this?

OpenStudy (anonymous):

is 15 - x in the denominator or just the 15?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the 15-x is the denominator

OpenStudy (anonymous):

20 5 ------ = ---- 15 - x x Anytime you have a fraction = a fraction, you can cross multiply. 20(x) = 5(15 - x) distribute 20x = 75 - 5x add 5x to both sides 25x = 75 divide by 25 x = 3

OpenStudy (anonymous):

thanks! will you help me out some more?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

140-x/81=x/9

OpenStudy (anonymous):

140 - x x ------- = --- 81 9 Because you have a fraction = a fraction - you cross multiply.

OpenStudy (anonymous):

(140 - x) * 9 = 81 * x

OpenStudy (anonymous):

now distribute

OpenStudy (anonymous):

1260 - 9x = 81x

OpenStudy (anonymous):

Add 9x to both sides

OpenStudy (anonymous):

1260 = 90x

OpenStudy (anonymous):

Divide both sides by 90

OpenStudy (anonymous):

14 = x

OpenStudy (anonymous):

thanks i get it now its a lot easier explained step by step like that you are a lot of help..